# Think-at-Hard: Selective Latent Iterations to Improve Reasoning Language Models

### v1

@@ Abstract
Improving reasoning capabilities of Large Language Models (LLMs), especially under parameter constraints, is crucial for real-world applications. Prior work proposes recurrent transformers, which allocate a fixed number of extra iterations per token to improve generation quality. After the first, standard forward pass, instead of verbalization, last-layer hidden states are fed back as inputs for additional iterations to refine token predictions. Yet we identify a latent overthinking phenomenon: easy token predictions that are already correct after the first pass are sometimes revised into errors in additional iterations. To address this, we propose Think-at-Hard (TaH), a dynamic latent thinking method that iterates deeper only at hard tokens. It employs a lightweight neural decider to trigger latent iterations, only at tokens that are likely incorrect after the standard forward pass. During latent iterations, Low-Rank Adaptation (LoRA) modules shift the LLM’s objective from general next-token prediction to focused hard-token refinement. We further introduce a duo-causal attention mechanism that extends attention from token sequence dimension to an additional iteration depth dimension. This enables cross-iteration information flow while maintaining full sequential parallelism. Experiments show that TaH boosts LLM reasoning performance across five challenging benchmarks while maintaining the same parameter count. Compared with baselines that iterate twice for all output tokens, TaH delivers 8.1-11.3% accuracy gains while exempting 94% of tokens from the second iteration. Against strong single-iteration Qwen3 models finetuned with the same data, it also delivers 4.0-5.0% accuracy gains. When allowing <<3% additional parameters from LoRA and iteration decider, the gains increase to 8.5-12.6% and 5.3-5.4%, respectively. Our code is available at https://github.com/thu-nics/TaH.
@@ 1 Introduction
Figure 1: Selective iteration can mitigate latent overthinking. (a) Toy example. Uniform latent iteration (always think-twice) can fix wrong predictions, but may also overthink and corrupt correct ones. (b) Next-token prediction accuracy of finetuned Qwen3-1.7B variants. Always think-twice causes more errors than corrections over direct reply. In contrast, the think-at-hard oracle, which iterates only when the first-pass prediction is wrong, achieves substantial improvements with minimal harm. While this oracle signal is unavailable in practice, it highlights the potential of selective iteration.
@@
Recent advances in Large Language Model (LLM) reasoning have enabled broad applications across diverse domains (jaech2024openai-o1; guo2025deepseek-r1; yang2025qwen3). With tens to hundreds of billions of parameters, LLMs can generate complex Chain-of-Thought (CoT) to solve challenging tasks. At the same time, smaller language models have also drawn increasing attention. With only a few billion parameters, they offer compelling alternatives: lower costs, faster inference, and suitability for edge computing (abdin2024phi3; allal2025smollm2; minicpm4_2025).
@@
At this crossroad, enhancing reasoning capabilities under parameter constraints becomes a central challenge. A common approach is to distill smaller models to mimic LLM CoT trajectories using next-token prediction supervision. However, not all tokens are equally predictable: certain tokens encode critical logic or reasoning directions that are fundamentally harder to predict (lin2024rho; fu2025r2r; wang2025beyond). With limited computation per output token, small models quickly hit a performance ceiling and mispredict some of these tokens. Once critical errors occur, the reasoning trajectory can irrecoverably diverge and produce drastically different outcomes.
@@
Prior work proposes recurrent transformers to address this parameter–performance paradox (hutchins2022block; saunshi2025loopedTrans; zeng2025pondering). Instead of verbalizing the next token immediately after one forward pass, these models typically feed the last-layer hidden states back into the LLM for additional passes in the latent space. Each pass refines the hidden representation without producing tokens. After a fixed number of iteration depths, the final hidden states are passed to the language modeling head to generate the next token. By uniformly allocating extra iterations per token, these models increase inference depth without enlarging parameter count, potentially benefiting hard reasoning tokens.
However, we identify a latent overthinking problem in fixed-depth recurrent transformers, where excessive iterations revise correct answers into wrong ones. As shown in Figure 1, finetuning Qwen3-1.7B-Base to always perform two iterations per token yields even more errors than the single-iteration baseline on the Open-R1 dataset (openr1). This occurs because most tokens are already predicted correctly in the first iteration, such as coherence or suffix tokens. Similar to overthinking in explicit CoT reasoning (wu2025morelesscot), latent overthinking on these easy tokens degrades performance despite extra computation. While the opposite latent underthinking exists for tokens that need more iterations to correctly predict, such cases are rarer. We define tokens that cannot be accurately predicted in a single forward pass as hard tokens, and ask our central question:
Can LLMs selectively dedicate latent iterations only to hard tokens?
If achieved, different iterations could specialize in distinct prediction focuses for more effective latent reasoning. Oracle experiments validate this approach: as shown in Table 4, a think-at-hard oracle improves MATH accuracy by 25-28%.
@@
Achieving dynamic latent iteration presents three main challenges. First, the model architecture should enable cross-depth attention, allowing each iteration to access full context. This is crucial because when early tokens skip deeper iterations, later tokens must still access their representations from shallower depths. Meanwhile, this cross-depth flow cannot compromise the sequence-level parallelism essential for efficient training and prefilling. Second, the model must adapt to changing objectives and distributions across iterations, while maximizing parameter reuse. Third, training must remain stable despite tight coupling dependencies: the iteration policy depends on prediction quality at each depth, while that quality depends on which tokens the policy sends to each depth.
@@
To address these challenges, we propose TaH, a dynamic latent thinking method that selectively applies deeper iterations only to hard tokens. As shown in Figure 2, TaH employs a neural decider to determine whether to continue iterating or verbalize each token. We design a duo-causal attention mechanism to enable cross-depth attention and full sequence parallelism. To specialize deeper iterations for hard-token refinement and preserve strong first-pass predictions, we apply LoRA adapters solely at iterations $d>1$. TaH is stably trained by aligning both LLM backbone and iteration decider with a static oracle iteration policy. We summarize our contributions as follows.
  • Selective Latent Iteration. We identify the latent overthinking phenomenon, revealing how false corrections harm easy tokens at redundant iterations. This insight guides our new paradigm where latent iteration depth adapts to token difficulty.
@@
  • Specialized Model Architecture. We develop a model architecture that natively supports selective iteration depths. The dedicated duo-causal attention mechanism, LoRA adapters, and iteration decider enable efficient cross-depth information flow, objective transitions, and dynamic depth selection.
@@
Experiments show that TaH consistently improves reasoning performance. Finetuned from Qwen3-0.6B-Base and 1.7B-Base with aligned parameter count, TaH achieves an average accuracy gain of 4.0-5.0% over standard single-iteration variants across five reasoning benchmarks, while applying deeper thinking to only 6% of tokens. With less than 3% additional parameters, these gains further increase to 5.3-5.4%. Compared with AlwaysThink which applies two iterations to all tokens, the gains are 8.1-11.3% and 8.5-12.6%, validating TaH’s high effectiveness.
@@ 2 Related Work
Unlike standard LLMs that verbalize at every autoregressive step, latent thinking shifts part of generation away from explicit natural-language CoT in order to improve reasoning (li2025implicit).
Signal-guided Control. These methods keep reasoning in token space but steers computation by inserting control tokens. Early work shows that simple filler tokens (e.g., dots) can mimic some benefits of CoT (pfau2404let). Building on this, later work expands the LLM vocabulary with [PAUSE] tokens and learns where to insert them for extra compute before predicting the next token (goyal2310think; kim2025learning). They are lightweight and easily integrable, but constrained to the discrete-token interventions with limited latent controls.
Latent Optimization. These methods perform autoregressive reasoning directly in internal representations, emitting little or no intermediate text. They distill and compress CoT into latent continuous embeddings through various strategies. Coconut and CCoT progressively replace text with latent thinking under final response supervision (hao2024training; cheng2024compressed); Token assorted and HCoT compress CoT spans to embeddings with hidden-state alignment (su2025token; liu2024expediting). SoftThink directly applies logit-weighted embeddings for latent iterations (zhang2025soft). While offering efficiency gains and flexible control over hidden trajectories, these methods sacrifice reasoning interpretability, with training-based ones further requiring heavy mitigation from strong verbal LLMs.
Recurrent Transformers. These methods interleave latent and verbal reasoning, introducing latent iterations before each token verbalization. After a standard forward pass, these methods feed latent states back as next-iteration inputs for a fixed number of iterations, then verbalize the output token. Existing approaches differ in the formation of next-iteration input. For example, Looped Transformer reuses last-layer hidden states directly (saunshi2025loopedTrans; geiping2025scaling), whereas Ponder uses logit-weighted embeddings (Equation 4(zeng2025pondering). Recurrent transformers combine advantages of visible reasoning trajectories with latent exploration. By reusing the parameters across iterations, it achieves deeper computation per token without parameter increases. However, the fixed depth burdens each iteration with both easy and hard tokens, potentially causing false corrections for already-correct predictions.
Positioning. TaH belongs to the recurrent transformer family but extends this paradigm significantly. It selectively allocates latent iterations to refine hard tokens, improving reasoning quality with specialized objectives across iterations. While concurrent works (bae2025mor; zhu2025ouro) also enable selective recursion, they require complete model retraining. TaH instead leverages existing pre-trained models, adding depth-aware LoRA and duo-causal attention to improve reasoning with minimal finetuning overhead.
3 Preliminary
Autoregressive LLMs. Modern LLMs generate text through an autoregressive next-token prediction process. It includes a prefill stage and a decode stage (radford2018gpt1; radford2019gpt2; kwon2023vllm). In the prefill stage, the model processes the entire input sequence in parallel; in the decode stage, it consumes one new token at a time along with cached history to predict the next token.
@@ 3 Preliminary
where $\mathcal{S}$ is a sampling rule such as greedy or nucleus sampling. Decoding repeats until an end-of-sequence token is generated.
Causal Attention. To respect autoregression, modern LLMs apply causal attention. As shown in Figure 2(a), each position attends only to itself and earlier positions, consistent with Equation 1. This design brings two key benefits: (1) it enables parallel training with next-token prediction and shifted logits, avoiding the need for token-by-token generation; and (2) it allows efficient inference by caching Key/Value states of past tokens instead of recomputing them.
Recurrent Transformers. Recurrent transformers introduce an inner loop that iterates in latent space before verbalizing each output token. Let $d\in\{1,2,\dots\}$ denote the iteration depth (written as a superscript), and set $x_{i}^{(0)}=E[t_{i}]$. At each iteration, recurrent transformers update $y_{i}$ with causal attention on the hidden states of the current iteration:
@@
An inner transition then produces the next-depth embedding. For example, Loop (saunshi2025loopedTrans) simply sets $x_{i}^{(d+1)}=y_{i}^{(d)}$, while Ponder (zeng2025pondering) uses a logit-weighted embeddings:
@@ 4.1 Duo-Causal Attention
Motivation. In recurrent transformers, attention typically operates within each iteration. For fixed-depth methods, standard causal attention on the current iteration’s Key and Value states already incorporates all context (Equation 3). However, dynamic iteration depths pose a challenge: tokens iterating at a deeper level cannot access the hidden states of previous tokens that verbalized at shallower depths. This creates a dilemma. On one hand, tokens require up-to-date states of all previous tokens for rich semantic context. On the other hand, efficient training requires all tokens at depth $d$ be computable in parallel, without depending on previous tokens’ deeper states ($d^{\prime}>d$) that have not yet been computed. Existing approaches compromise on one of these aspects. Some sacrifice parallelism by allowing attention to deeper iterations, forcing sequential generation during training (hao2024training); others preserve parallelism by restricting attention to only the initial iteration’s KVs (bae2025mor). To resolve this dilemma, we introduce a simple yet effective mechanism to maximize cross-depth information flow while maintaining high parallelism.
@@
When all tokens iterate only once (as in standard transformers), this naturally reduces to regular causal attention. The duo-causal design achieves both full parallel training and cross-depth information flow. At depth $d$, all tokens compute their depth-$d$ representations simultaneously using only and all information from depths $1$ through $d$. Moreover, duo-causal attention is fully compatible with existing systems, like FlashAttention (dao2022flashattention; dao2023flashattention2; shah2024flashattention3), requiring only simple modifications to the attention mask and KV cache (details in Appendix A.3.1).
4.2 Model Architecture
Motivation. Previous fixed-depth recurrent transformers use identical weights across all iterations. However, we find that over 85% of next-tokens are correctly predicted at the first iteration (Figure 1(b)) This suggests deeper iterations serve a different objective: they refine the first iteration’s prediction rather than predicting further ahead to the next-next token. This mirrors deep LLMs, where shallow layers predict next tokens for deeper layers to refine (belrose2023tunedLens; schuster2022confidentEarlyExit; bae2023fastEarlyExit). While deep LLMs naturally handle this shift through distinct parameters per depth, recurrent transformers must accommodate both objectives with shared weights, potentially limiting performance. Moreover, fixed iteration depths can cause latent overthinking, motivating our dynamic approach.
Backbone Model. To address the objective shift, we apply a LoRA adapter (hu2022lora) to the shared LLM backbone only for iterations $d>1$. As shown in Figure 2(c), this allows the base LLM to focus on latent embeddings, while the adapter handles the objective shift. It preserves strong next-token prediction at $d=1$, alleviating interference from deeper iterations. We also add residual connections across iterations to simplify the refinement and improve gradient flow. Formally, at depth $d$, we compute

### v2

@@ Abstract
Improving reasoning abilities of Large Language Models (LLMs), especially under parameter constraints, is crucial for real-world applications. Looped transformers address this by performing multiple latent iterations to refine each token beyond a single forward pass. However, we identify a latent overthinking phenomenon: most token predictions are already correct after the first pass, but are sometimes revised into errors in later iterations. In this work, we ask whether selectively skipping latent iterations may improve accuracy. We reveal significant potential with an oracle iteration policy that boosts model performance by up to 7.3%. Motivated by this, we propose Think-at-Hard (TaH), a looped transformer optimized for selective iteration. TaH employs a lightweight neural decider to trigger latent iteration only at tokens that are likely incorrect after the standard forward pass. During latent iterations, depth-aware Low-Rank Adaptation (LoRA) modules shift the LLM’s objective from general next-token prediction to focused hard-token refinement. A duo-causal attention mechanism extends attention from the token sequence dimension to an additional iteration depth dimension, enabling cross-iteration information flow with full sequential parallelism. Experiments on nine benchmarks show consistent gains across math, QA, and coding tasks. With identical parameter counts, TaH outperforms always-iterate baselines by 3.8–4.4% while skipping iterations on 93% of tokens, and exceeds single-iteration Qwen3 baselines by 3.0–3.8%. When allowing $3\%$ more parameters from LoRA and decider modules, the gains further increase to 5.3–6.2% and 6.1–6.8%, respectively. Our code is available at https://github.com/thu-nics/TaH.
@@ 1 Introduction
Figure 1: Latent iterations can fix wrong predictions, but can also overthink and flip correct ones. Selective iteration only when needed can improve reasoning with reduced computation.
@@
Recent advances in Large Language Model (LLM) reasoning have enabled broad applications across diverse domains (Jaech et al., 2024; Guo et al., 2025; Yang et al., 2025). With hundreds of billions of parameters, LLMs can generate complex Chain-of-Thought (CoT) to solve challenging tasks. At the same time, smaller language models have also drawn increasing attention. With only a few billion parameters, they offer compelling alternatives: lower costs, faster inference, and suitability for edge computing (Abdin et al., 2024; Team et al., 2025; Wang et al., 2025a).
@@
At this crossroads, enhancing reasoning capabilities under parameter constraints becomes a central challenge. A common approach is to distill smaller models to mimic LLM CoT trajectories. However, not all tokens are equally predictable: certain tokens encode critical logic or reasoning directions that are fundamentally harder to predict (Lin et al., 2024; Fu et al., 2025; Wang et al., 2025b). With limited computation per output token, small models quickly hit a performance ceiling and mispredict some of these tokens. Once critical errors occur, the reasoning trajectory can irrecoverably diverge and produce drastically different outcomes.
@@
Prior work proposes looped transformers to address this parameter–performance paradox (Hutchins et al., 2022; Saunshi et al., 2025; Zeng et al., 2025; Zhu et al., 2025). Instead of verbalizing the next token immediately after one forward pass, these models typically feed the last-layer hidden states back into the LLM for additional passes, refining representations in the latent space. After certain iterations, the final hidden states pass through the language modeling head to generate the next token. By uniformly scaling up iterations per token, these models can correct initially wrong token predictions, potentially increasing performance without increasing parameter count.
However, we identify a latent overthinking problem in looped transformers, where excessive iterations revise correct answers into wrong ones. As shown in Figure 3, while the second iteration corrects 8.7% of predictions, it also flips 2.1% of correct ones into errors. This occurs because most tokens, such as coherence or suffix tokens, are already predicted correctly after the first pass; further iterations may instead introduce harmful changes. This mirrors overthinking in explicit CoT reasoning (Wu et al., 2025), where additional reasoning steps degrade rather than improve answers. This reveals a surprising opportunity:
selectively skipping latent iterations on most tokens can further increase model performance.
We validate this with an oracle policy that iterates only on initially mispredicted tokens, as shown in Table 1. Compared to always iterating, this selective oracle can achieve up to 32% higher accuracy with optimized architecture.
@@
Achieving selective latent iteration presents three main challenges. First, the model architecture should enable cross-depth attention, allowing each iteration to access full context. This is crucial because when early tokens skip deeper iterations, later tokens must still access their representations from shallower depths. Meanwhile, this cross-depth flow cannot compromise the sequence-level parallelism essential for efficient training and prefilling. Second, the model must adapt to distribution shifts across iterations, while maximizing parameter reuse. Third, training must remain stable despite tightly coupled dependencies: the iteration policy depends on prediction quality at each depth, while that quality depends on the depths to which the policy assigns previous tokens.
Table 1: Performance comparison across iteration strategies, using Ouro-1.7B model except for the last row. Always1 means no latent iteration. Oracle policy iterates on 12–19% tokens. MMLU100 denotes the first 100 questions from MMLU-STEM.
Iter. Policy NTP AMC23 MMLU100 HE++
Always1 73.1 38.1 56.0 39.6
Always2 79.7 40.6 60.0 40.9
Orcl. 81.8/+2.1/+2.1 47.9/+7.3/+7.3 62.0/+2.0/+2.0 43.3/+2.4/+2.4
Orcl. w.TaH 89.3/+9.6/+9.6 68.8/+28.2/+28.2 85.0/+25.0/+25.0 72.9/+32.0/+32.0
@@
To address these challenges, we propose TaH, a looped transformer optimized for selective latent iteration. As shown in Figure 2, TaH employs a neural decider to determine whether to iterate or verbalize each token. We design a duo-causal attention mechanism to enable cross-depth attention with full sequence parallelism. To specialize deeper iterations for current-token refinement and preserve strong first-pass predictions, we apply LoRA adapters solely at iterations $d>1$. We train TaH stably by aligning both the LLM backbone and iteration decider with a static oracle iteration policy. We summarize our contributions as follows.
  • Selective Latent Iteration. We identify the latent overthinking phenomenon and quantify its influence on token prediction accuracy and downstream tasks. This insight motivates new directions where latent iteration is applied selectively to a few tokens for both better reasoning quality and efficiency.
@@
  • Specialized Model Architecture. We develop a model architecture that natively supports selective iterations. The dedicated duo-causal attention mechanism, LoRA adapters, and iteration decider enable efficient cross-depth information flow, objective transitions, and dynamic depth selection.
@@
We fine-tune TaH from Qwen3-Base 0.6B and 1.7B on Open-R1, then tested across nine reasoning benchmarks spanning math, QA, and coding tasks. TaH achieves average accuracy gains of 3.0% and 3.8% over standard single-iteration baselines. Compared to Ouro looped-transformer trained on the same data, TaH achieves 3.8–4.4% gains while reducing latent iterations by 93%.
@@ 2 Related Work
Unlike standard LLMs that verbalize at every autoregressive step, latent thinking shifts part of generation away from explicit natural-language CoT in order to improve reasoning (Li et al., 2025). We provide additional discussions on signal-guided control and latent optimization in Appendix A.5.
Looped Transformers. These methods interleave latent and verbal reasoning, adding latent iterations before each token verbalization. Previous work focuses on scaling up iteration depths, with the main architectural focus on next-iteration inputs: reusing hidden states directly (Saunshi et al., 2025; Geiping et al., 2025; Zhu et al., 2025) or using logit-weighted embeddings (Zeng et al., 2025). Looped transformers achieve deeper computation without parameter increases. However, uniform depth scaling burdens training and inference, and risks overthinking already-correct tokens.
Positioning. TaH belongs to the looped transformer family, but identifies selective iteration as a new design principle to improve performance. While concurrent works (Bae et al., 2025; Zhu et al., 2025) also enable dynamic recursion, they degrade performance at non-maximum depths and require full retraining. TaH instead leverages existing pre-trained models, adding depth-aware LoRA and duo-causal attention to improve reasoning with minimal fine-tuning overhead.
3 Preliminaries
Autoregressive LLMs. Modern LLMs generate text through an autoregressive next-token prediction process. It includes a prefill stage and a decode stage (Radford et al., 2018; 2019; Kwon et al., 2023). In the prefill stage, the model processes the entire input sequence in parallel; in the decode stage, it consumes one new token at a time along with cached history to predict the next token.
@@ 3 Preliminaries
where $\mathcal{S}$ is a sampling rule such as nucleus sampling. Decoding repeats until an end-of-sequence token is generated.
Causal Attention. Modern LLMs typically adopt causal attention mechanism. As shown in Figure 2(a), each position attends only to itself and earlier positions, consistent with Equation 1. This design brings two key benefits: (1) it enables parallel training with next-token prediction and shifted logits, avoiding the need for token-by-token generation; and (2) it allows efficient inference by caching Key/Value states of past tokens instead of recomputing them.
Looped Transformers. Looped transformers introduce an inner loop that iterates in latent space before verbalizing each output token. Let $d\in\{1,2,\dots\}$ denote the iteration depth (written as a superscript), and set $x_{i}^{(0)}=E[t_{i}]$. At each iteration, looped transformers update $y_{i}$ with causal attention on the hidden states of the current iteration:
@@
An inner transition then produces the next-depth embedding. For example, Loop (Saunshi et al., 2025) simply sets $x_{i}^{(d+1)}=y_{i}^{(d)}$, while Ponder (Zeng et al., 2025) uses a logit-weighted embeddings:
@@
4 Selective Latent Iteration Oracles
Setup. We analyze the latent iteration behavior of the Ouro looped transformer (Zhu et al., 2025) and our proposed TaH model (architecture detailed in Section 5). All models are fine-tuned from Qwen3-1.7B-Base on a balanced 100K-sample subset of the Open-R1 dataset (Hugging Face, 2025), following setups in Section 6.1.
Oracle Iteration Policy. To investigate the potential of selective iteration, we establish an oracle policy $\pi$. It triggers additional iterations only when the reference LLM $\theta$ mispredicts the target token at the first forward pass. In this section, we set $\theta$ to the model under evaluation, so $\pi$ acts as a greedy, locally optimal policy.
Formally, let $\hat{p}_{i}$ denote the reference model’s first-pass next-token distribution at position $i$, and $t_{i+1}$ the ground-truth token. The oracle iteration depth $d_{i}^{\pi}$ is:
$d_{i}^{\pi}\;=\;1+\mathcal{D}(\hat{p}_{i}^{(1)},t_{i+1}),$ (5)
where $\mathcal{D}$ is a binary discrepancy metric. We use top-1 mismatch here: $\mathcal{D}=\bm{1}[\hat{t}_{i+1}\neq t_{i+1}]$, with $\hat{t}_{i+1}=\arg\max_{t}\hat{p}_{i}^{(1)}$ ($\bm{1}[\cdot]$ denotes the indicator function). For simplicity, we assume $d_{\max}=2$. The arbitrary-depth case is in Appendix A.2.5, and alternative discrepancy metrics are ablated in Table 13.
For brevity, we call a token easy if the reference model correctly predicts it at the first pass ($d_{i}^{\pi}=1$), and hard otherwise. Prior work (Fu et al., 2025) shows that hard tokens typically occupy only a small proportion (e.g., $7\%$) of all tokens.
Next-Token Prediction. We evaluate next-token prediction (NTP) accuracy by comparing top-1 predictions against ground-truth tokens in the Open-R1 validation set. As shown in Table 1, always iterating twice improves NTP accuracy for Ouro by 6.6%, but at the cost of doubled depth. Surprisingly, the oracle policy, which skips iterations on 81-88% tokens, further improves accuracy by 2.1%. This gain comes from avoiding latent overthinking: without selective iteration, the model can revise correct predictions to wrong ones at the second pass, as shown in Figure 3.
Downstream Tasks. We next examine whether NTP gains translate to downstream task performance. Since ground-truth tokens are unavailable during generation, we use top-1 predictions from the stronger Qwen3-8B as proxy labels. Table 1 shows that the oracle policy improves downstream performance by 2.0–7.3%.
TaH Design Objectives. The oracle experiments reveal two key insights. First, models have significant untapped potential when iteration depth is selected correctly. Second, the oracle policy is effective enough to learn from, even though other globally optimal policies may exist. These findings motivate the design objectives of TaH: (1) its model architecture should natively support selective iteration depths; and (2) following the oracle policy, its training objective for deeper iterations is not to cover all tokens, but to selectively focus only on the few failing tokens. We later validate that TaH-1.7B can better utilize the oracle policy to achieve $25\%$ improvement, surpassing even Qwen3-4B. While the oracle requires ground-truth tokens unavailable at inference, approximating it with neural networks is a promising direction.
@@ 5.1 Duo-Causal Attention
Motivation. In looped transformers with fixed depth, standard causal attention on the current iteration’s KV states incorporates all context (Equation 3). However, dynamic iteration depths create a challenge: tokens at deeper levels cannot access hidden states of previous tokens that verbalized at shallower depths. This poses a dilemma between requiring up-to-date context from all previous tokens and maintaining parallel training where depth-$d$ computations cannot depend on uncomputed deeper states ($d^{\prime}>d$). Existing approaches compromise on one of these aspects. Some sacrifice parallelism by allowing attention to deeper iterations (Hao et al., 2024); others preserve parallelism by restricting attention to only the initial iteration’s KVs (Bae et al., 2025). To resolve this dilemma, we introduce a simple yet effective mechanism to maximize cross-depth information flow while maintaining high parallelism.
@@
When all tokens iterate only once (as in standard transformers), this reduces to regular causal attention. The duo-causal design achieves both full parallel training and cross-depth information flow. At depth $d$, all tokens compute their depth-$d$ representations simultaneously using only and all information from depths $1$ through $d$.
5.2 Depth-adaptive Model Architecture
Motivation. Previous looped transformers typically use identical weights across all iterations. However, we find that over 73% of next-tokens are correctly predicted at the first iteration (Table 1). This suggests deeper iterations serve a different objective: they refine the first iteration’s prediction rather than predicting further ahead to the next-next token. This mirrors deep LLMs, where shallow layers predict next tokens for deeper layers to refine (Belrose et al., 2023; Schuster et al., 2022; Bae et al., 2023). While deep LLMs naturally handle this shift through distinct parameters per depth, looped transformers must accommodate both objectives with shared weights, potentially limiting performance. Moreover, fixed iteration depths can cause latent overthinking, motivating our dynamic approach.
Backbone Model. To address the objective shift, we apply a LoRA adapter (Hu et al., 2022) to the shared LLM backbone only for iterations $d>1$. As shown in Figure 2(c), this allows the base LLM to focus on latent embeddings, while the adapter handles the objective shift. It preserves strong next-token prediction at $d=1$, alleviating interference from deeper iterations. We also add residual connections across iterations to simplify the refinement and improve gradient flow. Formally, at depth $d$, we compute